accents: é L^{A}TEX now inline: blaé test now inline: bla étest now inline: bla é test real spaces xxx tildes accent õ horizontal 2 - is – horizontal 3 - is — ligatures: ffl fi fl Yp what about vertical bars — horizontal bar - and tildes ¡ funny carrots ¿ copyright \begin{matrix}c \\ ◯ \end{matrix} pounds £ euro diamond◇ scandinavian ae æAE Æ o ø O Ø AA Å aa å Z ž alphaα eta η nu ν tau τ beta β theta θ xi ξ upsilon υ gamma γ iota ι phi ϕ delta δ kappa κ pi π chi χ epsilon ϵ lambda λ rho ρ psi ψ zeta ζ mu μ sigma σ

(omega ω) / (1 2 3)

4 5 6
7 8 9

so far not in verbatim (tt font)

now in verbatim (tt font)
accents: \'{e}
\LaTeX
now inline: bla\'{e} test
now inline: bla \'{e}test
now inline: bla \'{e} test
real spaces~~~xxx
tildes  accent \~o
horizontal 2 -  is --
horizontal 3 -  is ---
ligatures: ffl fi fl Yp
what about vertical bars |
horizontal bar -
and tildes ~
< funny carrots >
or other symbols: \\ { } [ ] _^? = ; / . , + * ( ) ' & % $ # " !